OpenStudy (anonymous):

If a+b=3 and ab=4, 1/a + 1/b = ? An explanation would also be helpful. Thanks!

OpenStudy (anonymous):

a+b/ab =3/4

jimthompson5910 (jim_thompson5910):

\[\frac{1}{a}+\frac{1}{b}\] \[\frac{b}{ab}+\frac{a}{ab}\] \[\frac{b+a}{ab}\] \[\frac{a+b}{ab}\] \[\frac{3}{4}\] So \[\frac{1}{a}+\frac{1}{b}=\frac{3}{4}\]

OpenStudy (anonymous):

Thank you! :D